OpenStudy (anonymous):

find decimal notation: 75 3/8%

OpenStudy (anonymous):

75.375

OpenStudy (anonymous):

Did you understand how to get the answer?

OpenStudy (anonymous):

\[75\frac{3}{8}=75+\frac{3}{8}=75+0.375\]

OpenStudy (anonymous):

I'm stil confused.. but somehow i stil got it

OpenStudy (phi):

first change 75 3/8 to a decimal. to change 3/8 into a decimal, divide 8 into 3  to get 75.375 but notice it was 75 3/8 % so this is 75.375 % to change the percent to a decimal, divide by 100 and drop the % sign: \[ \frac{75.375}{100}= 0.75375\] to divide by 100, move the decimal point left 2 places.